A. mass of 3.00 kg of water at 50 degree C has 1.0
Answer #1

mass of water, m = 3 kg
mass of ice, mi = 1 kg
let equilirium temp be T

a) Q gain = mCw(50 - T)
Q loss = mi*Lf + mi*Cw(T - 50)
Q gain = Q loss
mCw(50 - T) = mi*Lf + Mi*Cw(T - 50)
b) T = (mCw*50 - mi*Lf + Mi*Cw*50)/(mCw + MiCw)
c) T = (3*4186*50 - 1*335000 + 1*4186*50)/(3*4186 + 1*4186) = 29.99 deg